OpenStudy (anonymous):

find an equation at the tangent line to the curve y=e^x/(1+x^2) at the point (1,e/2).

OpenStudy (uri):

you differerentiate the curive first for the tangent of the line..then plug in \(x\) for the gradient.. show me your work first.

OpenStudy (anonymous):

y=e^x/(1+x^2) f'(x)=1+x^2(1)-e^x(x)/(1+x^2)^2 f'(1)=1+(1)^2-e^1/1+((1)^2)^2 f'(1)=2-e^1/4

OpenStudy (anonymous):

i don't no whether this is correct or incorrect.please can i have the answer please

OpenStudy (anonymous):

You stuffed up the numerator when differentiating the curve/function.

OpenStudy (anonymous):

You didn't do the quotient rule properly in the numerator.